ten samples of a process measuring the number of returns per 100 receipts were taken for a local retail store. the number of returns were 10, 9, 11, 7,3,12,8, 4,6,11.

find the standard deviation of the sampling distribution approximated to the fourth decimal point. (use p-bar formula)
Answer #1

$$ \begin{aligned} &10+9+11+7+3+12+8+4+6+11=81 \\ &81 /\left(10^{*} 100\right)=0.081 \\ &\sqrt{[0.081 *(1-0.081)] / 100}=0.0272835115 \end{aligned} $$